# MistralBSM: Leveraging Mistral-7B for Vehicular Networks Misbehavior Detection

### v1

@@ Abstract
Vehicular networks are exposed to various threats resulting from malicious attacks. These threats compromise the security and reliability of communications among road users, thereby jeopardizing road and traffic safety. One of the main vectors of these attacks within vehicular networks is misbehaving vehicles. To address this challenge, we propose deploying a pretrained Large Language Model (LLM)-empowered Misbehavior Detection System (MDS) within an edge-cloud detection framework. Specifically, we fine-tune Mistral-7B, a state-of-the-art LLM, as the edge component to enable real-time detection, whereas a larger LLM deployed in the cloud can conduct a more comprehensive analysis. Our experiments conducted on the extended VeReMi dataset demonstrate Mistral-7B’s superior performance, achieving 98% accuracy compared to other LLMs such as LLAMA2-7B and RoBERTa. Additionally, we investigate the impact of window size on computational costs to optimize deployment efficiency. Leveraging LLMs in MDS shows interesting results in improving the detection of vehicle misbehavior, consequently strengthening vehicular network security to ensure the safety of road users.
Index Terms:
Vehicular network, Large Language Models (LLMs), misbehavior detection, intrusion detection
@@ I Introduction
The evolution of vehicular networks [1], commonly known as vehicle-to-everything (V2X), has been marked by continuous developments in communication and transportation technologies. Originally conceived as a means of improving vehicle safety through the exchange of information between vehicles and the infrastructure, research expanded into areas such as cooperative driving, advanced scenarios to support automated driving, and the integration of vehicle networks with emerging technologies such as 5G and beyond [2] and the Internet of Things (IoT) [3]. Today, as the potential of connected autonomous vehicles promises to redefine the future of private and public transportation, the evolution of vehicular networks research continues apace, particularly with regard to cybersecurity issues[4]
@@
Indeed, vehicular networks are exposed to various threats resulting from malicious attacks [5]. These threats compromise the security and reliability of communications among road users, thereby jeopardizing road and traffic safety. One of the main vectors of these attacks within vehicular networks is misbehaving vehicles [6]. Misbehaving vehicles can launch denial-of-service (DoS) attacks which flood the network with malicious traffic, disrupting communication between vehicles [7]. As a result, this disruption can lead to communication breakdowns, making vehicles unaware of each other and disrupting the operation of both traffic systems and vehicular safety applications. In addition, message tampering and integrity attacks can mislead on-board control systems, compromising vehicle safety. In this scenario, misbehaving vehicles deliberately transmit inaccurate or erroneous messages despite normal hardware and software functioning [8]. Vehicle misbehavior can result in significant repercussions, including serious traffic disruptions or even accidents. Hence, detecting misbehavior is crucial in vehicular communication systems. However, it is a complex challenge due to the dynamic nature of traffic environments and the diverse forms misbehavior can take. Vehicles interact unpredictably, leading to various behaviors that may differ from expected patterns. Additionally, the large volume of data generated by vehicles adds complexity. Analyzing this data in real-time requires advanced algorithms and computing resources.
@@
Machine learning (ML) has recently emerged as an interesting technology in the development of misbehavior detection systems (MDS). ML enables MDS to adapt and learn from historical data, continuously improving its ability to identify abnormal or malicious behavior. [9]. With the recent advancements in large language models (LLMs) there is an exciting opportunity to further bolster the capabilities of MDS. Indeed, models, such as the GPTs [10], LLAMA-2[11] and BERT[12] have demonstrated remarkable capabilities in various natural language processing tasks, but also showed interesting capacities in network security tasks [13], including for vehicular networks security [14, 15].
@@
The potential of LLMs in vehicle misbehavior detection stems from the fact that LLMs can facilitate the extraction of insights from textual data of vehicle communications. Through semantic analysis and contextual understanding, these models can identify linguistic features that can signify potential abnormal behavior patterns. However, using LLMs for vehicle misbehavior detection poses significant challenges, notably in privacy and resource constraints. Given their extensive parameter counts, often in the tens of billions or even trillions, LLMs are naturally suited for cloud deployment. Nevertheless, transferring all vehicle communication data to cloud infrastructure poses potential risks for real-time misbehavior detection and raises concerns about data privacy.
@@
Recently, a new large-scale model named Mistal-7B [16] has been introduced, rivaling existing models with fewer parameters thanks to various architectural choices. This opens the possibility to use an LLM-empowered MDS architecture with edge-cloud collaboration, where a light LLM Model such as Mistral-7B resides at the edge (e.g., at a roadside unit), and a bigger LLM resides in the cloud. Edge LLMs can perform initial vehicle misbehavior detection in real-time, providing immediate feedback and alerts to vehicle systems. Cloud LLMs can complement this by conducting deeper analysis, incorporating additional contextual information from a broader, and possibly multi-modal dataset stored in the cloud (e.g., multimodal traffic information from the intelligent transportation system). Using the edge LLM has the advantage of processing vehicular communication data locally, minimizing the need to transmit raw data to the cloud and preserving user privacy. Cloud LLMs can be utilized for more general analysis that uses edge LLMs insights, and other data external to the communication system, ensuring compliance with privacy regulations and minimizing the risk of data privacy breaches.
@@
In this article, we explore the use of Mistral-7B as an edge LLM to be located at a roadside unit, and we explore the capacity of the LLM to classify each vehicle as normal or misbehaving. The main contributions of this article are as follows:
  • We use Mistral-7B for misbehavior detection based on the analysis of observed basic safety massages (BSMs).
  • We transpose a vehicular misbehavior detection task into a sequence classification by converting windows of BSMs originating from a specific vehicle into text prompts to be classified by a language model.
  • We fine-tune Mistral-7B as a sequence classifier to assess a vehicle behavior nature through a sequence of its broadcasted BSMs, resulting in MistralBSM.
@@
  • We assess Mistral-7B’s performance compared to other prominent open-source LLMs, namely LLAMA2-7B and BERT based architecture, in the context of MDS, and we demonstrate its superior results.
@@
The rest of this paper is organized as follows. Section II introduces the related works of misbehavior detection methods in vehicular networks. Section III introduces a background about vehicular networks architecture, LLMs and low rank adaptation for LLM fine-tuning. In Section IV, we define our system model, the proposed solution and introduce Mistral for misbehavior detection. in Section V, we do a performance evaluation of our model and compare our results.
@@ II Related works
II-A Vehicular misbehavior detection
Detecting misbehavior is widely recognized as a critical task in V2X security. ML-based approaches are commonly favored due to their distinctive ability to model and classify misbehavior, surpassing other rule-based strategies.
In [17], various models including Random Forest, Decision Tree, Ada Boost Classifier, Ridge Classifier, Naive Bayes, and XGBoost have been evaluated on the CAN-OITDS dataset for multi-class classification involving four distinct attack types. The results showed an advantage of XGBoost. The authors of [18], also employed ML technique for misbehavior detection. The proposed framework enhances the detection performance by including movement and position plausibilities with the BSM features in the input of the classifier. The considered algorithm includes SVM, KNN, Naive Bayes, Random Forest, ensemble boosting and voting ensemble. The primary drawback of such models is their limited ability to adequately handle the temporal aspect of specific attacks aimed at V2X communication.
To address this gap, some researchers considered using LSTM based architectures. For example, in [19], [20] and [21], LSTM and CNN were combined to detect misbehavior, yielding promising results. On the other hand, in [22] and [23], considered the misbehavior detection as time-series anomaly detection problem that could be modeled by a Markov Decision process where each decision is based on the previous ones. They proposed a DQN-based classifier where the agent q-network is formed by LSTM layers and a fully connected layer. Particularly, in [22], the proposed framework uses KNN as an unsupervised algorithm for labelling BSMs. Thereafter, the labeled data is used by the DQN agent to learn the optimal policy. The effectiveness of these techniques underscores the importance of considering the temporal dimension in detecting misbehavior within vehicular networks, where LSTM has demonstrated efficacy. However, interest has shifted with the introduction of transformers and LLMs, as they enable efficient processing of sequenced data.
II-B LLMs for network security
LLMs have showcased remarkable capabilities that extend beyond natural language processing (NLP), generating considerable interest across diverse domains. Notably, they have gained interest in network security, with encoder-based transformers being particularly renowned.
For example, BERT served as the foundation for log anomaly detection systems as described in ([24], [13], [25]). In [24], BERT was trained on normal log key sequences extracted from raw network logs to predict the next masked key. Subsequently, during the detection phase, the model predicted normal keys with high precision, while abnormal keys exhibited low detection accuracy. Conversely, in [13], the authors fine-tuned BERT using a fully connected classifier, where the input consists of a time window comprising a sequence of parsed log data. In [25], log anomaly detection is achieved by fine-tuning LongFormer, a transformer based on BERT, exclusively using normal log data for masked token prediction. The threshold for discerning deviations in new log data from established normal patterns was determined based on the accuracy of top-k token prediction.
In the vehicular domain, a two-stage approach for detecting intrusions in the Controller Area Network (CAN) was proposed [15]. The approach involves pre-training BERT on normal CAN data to understand its semantic representation, followed by fine-tuning the model to classify messages as either normal or malicious. Despite its promising results, the model evaluates each packet individually, overlooking the sequential nature of some attacks. In contrast, in [26], a sliding window of CAN-IDs is fed to BERT to detect message injection in the CAN-bus. The model is trained using masked language self-supervised training to predict a set of candidates CAN IDs for a masked ID. If the actual ID does not correspond to one of the predicted IDs, it is considered abnormal. The results indicate that BERT-based approaches outperform other ML methods such as PCA and Autoencoders in terms of anomaly detection performance. In another work [27] the authors explore the use of well-known LLMs for CAN intrusion detection. They examine different architectures where an LLM is combined with a classification head, utilizing models like SecureBERT, BERT, and LLAMA-2. These models are fine-tuned on a database related to CAN. The results confirm the effectiveness of the models and highlight the superiority of larger models such as LLAMA-2 in capturing intricate message distributions.
In addition to research concentrating on in-vehicle communication security, a transformer was utilized for detecting misbehavior in V2X communication, as discussed in [14]. The authors fine-tuned an encoder-based transformer using normal BSM sequences, each with a length of 200, to replicate them. Subsequently, misbehavior detection was accomplished by comparing the differences between a misbehaving vehicle’s sequence and its transformed counterpart against a predefined threshold. While the results are promising, processing a 200 BSM input sequence may pose computational limitations, as can behavior evaluation with an insufficient number of BSMs. Another work using transformers is presented in [28], where a privacy-preserving communication architecture is proposed, utilizing a transformer-based Intrusion Detection System (IDS) trained on edge devices with data from connected vehicles. The cloud aggregates the model weights to distribute the final global model. The IDS utilizes a tailored encoder-based transformer, which integrates feature attention and routing, to choose the best feature sets from network packets for classification. Despite the promising results, a limitation of this work arises from the choice of a packet-wise approach, disregarding the importance of the packet context for the detection. In addition, conducting tests on non-vehicular data raises questions regarding the model’s effectiveness in detecting specific attacks on vehicular networks.
@@ III-A Vehicular communication models
V2X communications encompass a variety of technologies facilitating real-time information exchange between vehicles and their surroundings, thereby enhancing road safety, and traffic efficiency. Figure 1 provides an overview of the different types.
@@
In contrast, V2I communication allows vehicles to interact with roadside units (RSUs) such as streetlamps, RFID readers, lane markers, and signage. This bidirectional communication enables the real-time exchange of information about road conditions, traffic, accidents, and more, thereby enhancing overall traffic management and safety[29].
V2P, or vehicle-to-pedestrian communication, involves interactions with vulnerable entities in the environment, such as pedestrians and cyclists. These individuals typically carry smartphones or wearable devices capable of transmitting their location to alert nearby vehicles of their presence, contributing to enhanced safety for all road users.
Additionally, V2X includes V2N communication, which enables vehicles to connect to cloud-based services and platforms. This connectivity facilitates access to a wide range of services such as over-the-air updates, remote diagnostics, real-time traffic and navigation information, and data analytics for traffic management purposes.
In order to protect ensure vehicle privacy, vehicles use pseudonyms for communications. Vehicle pseudonyms are temporary, randomly generated identifiers used in vehicular communication systems to provide privacy and anonymity for vehicles while still allowing them to communicate with each other and infrastructure. These pseudonyms are periodically changed to prevent tracking of vehicles over time.
@@ III-B Transformers and LLMs
The transformer is a groundbreaking architecture that revolutionized the task of sequenced data processing, introducing innovative mechanisms for capturing complex dependencies and patterns within sequences. This was achieved by leveraging the self-attention mechanism to weight the importance of each token in sequence with regard to the other tokens. The weights are obtained by applying three linear transformations, namely (keys (k), queries (Q) and values (V)) on the input sequence of size $d_{k}$ following the equation 1.
Several variations of transformers have been proposed and customized to address a broad spectrum of tasks, encompassing language understanding, generation, translation, and more. Notable examples include the original Transformer architecture [30] which is presented in Figure 2, comprised of both encoder and decoder blocks, as well as BERT [12] and RoBERTa [31], both of which are encoder-only transformers. Additionally, LLAMA [11] and GPT exemplify the decoder-only type of transformers, primarily designed for generation tasks.
Mistral-7B, is also a decoder-only transformers that employs specific design decisions, allowing it to outperform models such as the 13 billion-parameter variation of Llama-2 and the 34 billion-parameter Llama-1 with only 7 billion parameters. Those technical choices involve using Sliding Window Attention (SWA) that operates by defining a fixed-size window surrounding each token, thus limiting the number of tokens it has to tend for [32]. Compared to the vanilla self-attention used in [30], this technique enables longer sequence processing with lower computation cost. Moreover, Mistral-7B employs the Grouped-query attention (GQA) technique where a group of queries shares a single key and value heads[33], thus speeding up the inference time of the model.

### v2

# MistralBSM: Leveraging The Mistral-7B LLM for Vehicular Networks Misbehavior Detection

@@ Abstract
Malicious attacks on vehicular networks pose a serious threat to road safety as well as communication reliability. A major source of these threats stems from misbehaving vehicles within the network. To address this challenge, we propose a Large Language Model (LLM)-empowered Misbehavior Detection System (MDS) within an edge-cloud detection framework. Specifically, we fine-tune Mistral-7B, a compact and high-performing LLM, to detect misbehavior based on Basic Safety Messages (BSM) sequences as the edge component for real-time detection, while a larger LLM deployed in the cloud validates and reinforces the edge model’s detection through a more comprehensive analysis. By updating only 0.012% of the model parameters, our model, which we named MistralBSM, achieves 98% accuracy in binary classification and 96% in multiclass classification on a selected set of attacks from VeReMi dataset, outperforming LLAMA2-7B and RoBERTa. Our results validate the potential of LLMs in MDS, showing a significant promise in strengthening vehicular network security to better ensure the safety of road users.
@@ I Introduction
The evolution of vehicular networks [1], commonly known as vehicle-to-everything (V2X), has been marked by continuous developments in communication, transportation, and consumer technologies. Originally conceived as a means of improving vehicle safety through the exchange of information between vehicles and the infrastructure, research expanded to include more advanced scenarios such as cooperative, automated driving and more consumer-centric applications such as in-vehicle infotainment systems, personalized mobility experiences, and seamless connectivity with consumer personal devices [2]. This evolution has been accelerated by the integration of emerging technologies like 5G and beyond [3] and the Internet of Things (IoT) [4], which enable more sophisticated, consumer-oriented services.
Today, as the potential of connected autonomous vehicles promises to reshape the landscape of personal and public transportation, the evolution of vehicular networks research continues apace, focusing on enhancing consumer experiences. This includes addressing network security challenges [5], enabling a reliable and secure experience within the intelligent transportation scene.
@@
Indeed, vehicular networks are exposed to various threats resulting from malicious attacks [6]. These threats compromise the security and reliability of communications among road users, thereby jeopardizing road and traffic safety. One of the main vectors of these attacks within vehicular networks is vehicle misbehavior [7]. Misbehaving vehicles can launch denial-of-service (DoS) attacks, which flood the network with malicious traffic, disrupting communication between vehicles [8]. As a result, this disruption can lead to communication breakdowns, making vehicles unaware of each other and disrupting the operation of both traffic systems and vehicular safety applications. In addition, message tampering and integrity attacks can mislead on-board control systems, compromising vehicle safety. In this scenario, misbehaving vehicles deliberately transmit inaccurate or erroneous messages despite normal hardware and software functioning [9]. Vehicle misbehavior can result in significant repercussions, including serious traffic disruptions or even accidents. Hence, detecting misbehavior is crucial in vehicular communication systems. However, it is a complex challenge due to the dynamic nature of traffic environments and the diverse forms misbehavior can take. Vehicles interact unpredictably, leading to various behaviors that may differ from expected patterns. Additionally, the large volume of data generated by vehicles increases the complexity. Analyzing this data in real-time requires advanced algorithms and computing resources.
@@
Machine learning (ML) has recently emerged as an interesting solution in the development of Misbehavior Detection Systems (MDS). ML enables MDS to adapt and learn from historical data, continuously improving its ability to identify abnormal or malicious behavior [10]. With the recent advancements in large language models (LLMs) there is an exciting opportunity to further bolster the capabilities of MDS. Indeed, models such as the GPTs [11], LLAMA-2[12] and BERT [13] have demonstrated remarkable capabilities in various natural language processing tasks but also showed interesting capacities in network security tasks [14, 15, 16], including for vehicular networks’ security [17, 18].
@@
The potential of LLMs in vehicle misbehavior detection stems from the fact that LLMs can facilitate the extraction of insights from textual data of vehicle communications. Through semantic analysis and contextual understanding, these models can identify relevent features that can signify potential abnormal behavior patterns. However, using LLMs for vehicle misbehavior detection poses significant challenges, notably in privacy and resource constraints. Given their extensive parameter counts, often in the tens of billions or even trillions, LLMs are naturally suited for cloud deployment. Nevertheless, transferring all vehicle communication data to cloud infrastructure poses potential risks for real-time misbehavior detection and raises concerns about data privacy.
@@
Recently, several LLMs were introduced, among which Mistral-7B [19] has gained attention for its exceptional balance between model size and performance, making it particularly suitable for ressource constrainted application like vehicular networks. More importantly, it is open-weight model, which opens the possibility to adapt and leverage it as an LLM-empowered MDS architecture. In this work we propose an edge-cloud setting, where a quantized Mistral-7B is deployed at the edge (e.g., at a roadside unit (RSU), and a deeper LLM resides in the cloud. The Edge LLM can perform initial vehicle misbehavior detection in real-time on short Basic Safety Message (BSM) sequences, providing immediate feedback and alerts to vehicle systems. Cloud LLMs can complement this by conducting more in-depth analysis on longer BSM sequences and by incorporating additional contextual information from a broader and possibly multi-modal dataset stored in the cloud (e.g., multimodal traffic information from the intelligent transportation system). Using the edge LLM has the advantage of processing vehicular communication data locally, minimizing the need to transmit raw data to the cloud and preserving user privacy. Cloud LLMs can be utilized for more general analysis that uses edge LLMs insights and other data external to the communication system, ensuring compliance with privacy regulations and minimizing the risk of data privacy breaches.
@@
In this article, we focus mainly on the implementation of Mistral-7B as the edge LLM to classify vehicle behaviors. Thus, the main contributions of this article are as follows:
  • We propose an LLM-based Misbehavior detection framework leveraging Mistral-7B for BSM sequence classification.
  • We design our data preprocessing carefully to avoid label leaking or training data leaking for a more reliable performance assessment and efficient model inference to enable easy edge deployment.
@@
  • We compare Mistral-7B’s performance to other prominent open-source LLMs, namely LLAMA2-7B and BERT-based architecture, in the context of MDS, and we demonstrate its superior results.
  • We use quantization and study BSM sequence lengths to optimize model deployment within the limited computational resources of vehicular networks.
@@
The rest of this paper is organized as follows. Section II introduces the related works of misbehavior detection methods in vehicular networks. Section III introduces a background about vehicular networks’ architecture, LLMs and low-rank adaptation for LLM fine-tuning. In Section IV, we define our system model and introduce Mistral for misbehavior detection. In Section V, we do a performance evaluation of our model and compare our results.
@@ II Related works
Misbehavior detection is a pivotal component in vehicular networks where ML can offer an advantage over rule-based approaches given the significant amount of communication data generated by the network and required to achieve an accurate detection. In the literature, several ML-approaches have been proposed, from simple algorithms to deep models, to address in-vehicle attacks as well as V2X communications misbehavior. For instance, the authors [20] studied several famous ML algorithms, including tree-based models such as Decision Tree, XGBoost, and Ada Boost classifiers in addition to Ridge classifier and Naive Bayes on CAN-OITDS dataset, where XGBoost achieved the best performance. On the other side, the authors of [21] studied ML for V2X anomaly detection using models such as Support Vector Machines (SVM), K-Nearest Neighbors (KNN), Naive Bayes, Random Forest, ensemble boosting, and voting ensemble, and were able to achieve good classification results. Despite being lightweight and achieving promising results, simple ML-techniques lack the ability to process the time facet of network data.
This drawback was addressed by LSTM-based architectures that were designed to effectively model temporal data. In  [22, 23, 24], the authors combined LSTM and CNN to derive spatial and temporal features, achieving good misbehavior detection. Additionally, the time aspect was further emphasized in [25] where the authors approached misbehavior detection as a time-series anomaly detection problem, modeled as a Markov Decision Process (MDP), where each decision depends on preceding states. The improvement introduced by these approaches highlights the importance of the temporal dimension in detecting misbehavior within vehicular networks.
The emergence of transformers shifted the interest due to their unprecedented capabilities to efficiently process sequenced data types. In vehicular networks, CAN-bus anomaly detection took advantage of these capabilities of transformer-based models and showed that BERT-inspired architectures surpass other deep models [18, 26, 27].Beyond works focusing exclusively on CAN-bus, the authors in [28] propose a hybrid Network IDS (NIDS), named IoV-BERT-IDS, for both in-vehicle (CAN-bus) and extra-vehicle networks. The model leverages a semantic extractor to convert raw CAN-Bus and network sequences into BERT-compatible representations, pretrained using Masked Byte Word Model (MBWM and Next Byte Sentence Prediction (NBSP) tasks to capture contextual and sequential features. It is then fine-tuned for intrusion detection, achieving about 99% accuracy across CICIDS, BOT-IOT, Car-Hacking, and IVN-IDS datasets. However, these datasets are not specifically V2X communication datasets; as a consequence, we cannot assume systematically that the model will give the same performance in our study case.
On the other hand, the authors in [17] focused on V2X communications misbehavior by fine-tuning an encoder-based vanilla transformer in a semi-supervised approach to detect anomalies. Their model reconstructs vehicle behavior by regenerating position and speed from sequences of 200 BSMs, then detects misbehavior based on the Mean Squared Error (MSE) between the original and reconstructed features compared to a threshold, yet their data preparation includes a hint to the vehicle’s ID in the dataset, which can lead to data label leakage. Additionally, Larger LLM models such as GPT-4o, LLaMA-3.2 and Gemini 1.5 Pro were used in [29] alongside DNNs such as VGG19 and ResNeXt to detect motion forging and natural denoising diffusion (NDD) attacks. The results indicate that GPT-4o delivers the strongest performance overall, though the model is not open source, which in a practical scenario will require sharing sensitive vehicular network telemetry with a third party, raising serious privacy and security concerns. Furthermore, LLMs were coupled with explainable AI technics explainable AI technics such Kappa coefficient, Matthews Correlation Coefficient, and SHapley Additive exPlanation (SHAP), in [30], to validate the model’s misbehavior detection capabilities. This approach studies aims to understand the reason behind a vehicle behavior classification achieved by LLM, however the SHAP results strongly suggest that the message ID feature has the biggest influence on the classification, raising potential overfitting or train/test data leakage.
From another perspective, the authors in [31] propose a transformer-based Intrusion Detection System (IDS) for V2X trained within federated learning approach. Their model is a tailored encoder-based transformer that integrates feature attention and routing to select the most suitable feature set for network packet classification. The proposed system overlooks temporal context of the packet in classification, which can be considered a major drawback. This limitation is addressed in [32], where the authors train their transformer encoder-based model to classify BSM-windows, showing that the federated learning gives a model with better classification accuracy compared to the standalone model.
Despite the growing use of LLMs for misbehavior detection, many studies provide limited details on data preparation. The results of some works strongly suggest potential training data leakage, particularly when identifiers such as message ID or pseudo-ID features are used, making it difficult to attribute performance gains solely to the model’s learning capacity. Moreover, it remains crucial to investigate the computational resources required effectively to deploy such models in real-world settings.
@@ III-A Vehicular Communication Models
V2X communications encompass various technologies facilitating real-time information exchange between vehicles and their surroundings, thereby enhancing road safety, and traffic efficiency. Fig.1 provides an overview of the different types.
@@
In contrast, V2I communication allows vehicles to interact with RSUs such as streetlamps, RFID readers, lane markers, and signage. This bidirectional communication enables the real-time exchange of information about road conditions, traffic, accidents, and more, thereby enhancing overall traffic management and safety[33].
V2P, or vehicle-to-pedestrian communication, involves interactions with Vulnerable Road Users (VRU), such as pedestrians, cyclists and motor wheelers [34]. These individuals typically carry smartphones or wearable devices capable of transmitting their location to alert nearby vehicles of their presence, contributing to enhanced safety for all road users.
Additionally, V2X includes V2N communication, which enables vehicles to connect to cloud-based services and platforms. This connectivity facilitates access to a wide range of services, such as over-the-air updates, remote diagnostics, real-time traffic and navigation information, and data analytics for traffic management purposes.
In order to protect ensure vehicle privacy, vehicles use pseudonyms for communications. Vehicle pseudonyms are temporary, randomly generated identifiers used in vehicular communication systems to provide privacy and anonymity for vehicles while still allowing them to communicate with each other and infrastructure. These pseudonyms are periodically changed to prevent tracking of vehicles over time [35].
@@ III-B Transformers and LLMs
The transformer is a groundbreaking architecture that revolutionized the task of sequenced data processing, introducing innovative mechanisms for capturing complex dependencies and patterns within sequences. This was achieved by leveraging the self-attention mechanism to weight the importance of each token in sequence compared to the other tokens. The weights are obtained by applying three linear transformations, namely (keys (k), queries (Q) and values (V)) on the input sequence of size $d_{k}$ following the equation 1.
$\text{Attention (Q, K,V)}=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
Several variations of transformers have been proposed and customized to address a broad spectrum of tasks, encompassing language understanding, generation, translation, and more. Notable examples include the original Transformer architecture [36] which is presented in Fig. 2, formed by both encoder and decoder blocks. BERT [13] and RoBERTa [37], are another type of architecture formed only by encoder blocks. On the other hand, LLAMA [12] and GPT are decoder-only types of transformers, primarily designed for generation tasks. The difference between the encoder-based and decoder-based transformer is self-attention module type. While in the case of BERT-like models, the attention is bidirectional, meaning that all the sequence tokens are used to compute the attention score for each token. The decoder-based models use masked self-attention technique, in which computing the attention score of a token requires only past tokens
Mistral-7B is another example of a decoder-based transformer that leverages specific design decisions, allowing it to outperform models such as the 13 billion-parameter variation of Llama-2 and the 34 billion-parameter Llama-1 with only 7 billion parameters. Those technical choices involve using Sliding Window Attention (SWA) that operates by defining a fixed-size window surrounding each token, thus limiting the number of tokens it has to tend to [38]. Compared to the vanilla self-attention used in [36], this technique enables longer sequence processing with lower computation cost. Moreover, it employs the Grouped-query attention (GQA) technique, where a group of queries shares a single key and value head [39], thus speeding up the model inference time.
Mistral-7B is suited for misbehavior detection in vehicular networks because it combines the performance of larger models with a more compact size. Its smaller size and efficient sequence processing make it more suitable for deployment at the edge, where computational resources are often limited [40]. Additionally, Mistral-7B’s pretraining on a large and diverse dataset enables easier fine-tuning for specific tasks like ours, requiring less data for adaptation. This advantage reduces the overall cost of supervised learning while leveraging its powerful ability to extract semantic information from sequences. Overall, Mistral-7B’s capabilities make it suitable for misbehavior detection in vehicular networks.
Figure 2: Transformer architecture
III-C Low-rank Adaptation
In traditional fine-tuning, the model’s weights are directly modified to adapt the model to the new considered task. While fine-tuning is a crucial approach in ensuring transfer of knowledge, the process becomes computationally daunting as the number of parameters increases. Notably, fine-tuning large language models poses significant challenges in terms of memory requirements and computational resources.